# The Computational Limits of State-Space Models and Mamba via the Lens of Circuit Complexity

### v1

@@
Yifang Chen yifangc@uchicago.edu. University of Chicago.    Xiaoyu Li xli216@stevens.edu. Stevens Institute of Technology.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
@@ 1 Introduction
Sequential neural networks like RNNs, including their variants such as LSTMs and GRUs [36, 15], have good performance in capturing temporal dependencies and processing input step-by-step [13]. These advantages make them effective in tasks including time-series prediction [4] and speech recognition [62]. Traditional RNNs [37] and their enhanced variance, LSTMs perform well in testing because of their sequential nature, but their training times tend to be slow and suffer from vanishing or exploding gradient issues, which limit their capabilities to capture long-term dependencies [73]. Transformers [66], equipped with a self-attention mechanism, provide an efficient solution to the slow training problem by enabling parallelized computations. Large Language Models (LLMs) based on the Transformer architecture, such as GPT-4 [58], GPT-4o [59], OpenAI’s o1 [60], Llama 3.1 [53], Claude [3], and Gemini [25], have become ubiquitous nowadays, and their integrations to modern technology reshaped our expectations of the limits of their capabilities. Transformers are capable of training efficiently on large datasets, but their quadratic memory and time complexity with respect to sequence length make them expensive in resources, both in terms of memory and processing power, during training and inference. Specifically, self-attention mechanisms grows $O(n^{2})$ in terms of computational complexity [44].
@@
In Section 2, introduce the works related our paper. Section 3 introduces key computational concepts and Mamba definitions that form the basis for subsequent sections. Then, we present the circuit complexity results for Selective SSM and Mamba in Section 4. Section 5 details our hardness results. Finally, Section 6 gives a conclusion.
@@ 2 Related Work
Circuit Complexity, a crucial set of metrics in computational complexity theory, studies the computational power of circuit families. It has valuable applications in comprehending the capabilities of machine learning models [61, 32, 31, 56, 55, 21, 47, 74, 19]. The complexity classes include $\mathsf{AC^{0}}$ represents problems that are highly parallelizable equipped with standard logic gates, which can be solved by constant-depth circuits with unbounded fan-in $\mathsf{AND}$, $\mathsf{OR}$, and $\mathsf{NOT}$ gates; $\mathsf{TC^{0}}$ class extends from $\mathsf{AC^{0}}$ with additional majority gates; $\mathsf{NC^{1}}$ problems can be solved by $O(\log n)$-depth circuits with bounded fan-in. These circuit complexity classes form a hierarchy: $\mathsf{AC}^{0}\subset\mathsf{TC}^{0}\subseteq\mathsf{NC}^{1}$ [56]. The question of whether $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$ remains an open topic of discussion. [41] demonstrates that while Transformers can simulate nonsolvable semi-automata, their depth is influenced by the length of the input sequence. Building on this, [47] investigates the expressive power of Transformers augmented with Chain-of-Thought (CoT) reasoning in the context of circuit complexity. They propose the following relationships:
@@
Transformers have shown outstanding performance on tasks from natural language processing, but they present limited effectiveness in mathematical computations. A series of research highlights the reasoning limitations of Transformer Model [2, 55, 12, 70, 48, 14][14] shows that average-hard attention transformers (AHATs) and softmax-attention transformers (SMATs) are in $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ with $O(\operatorname{poly}(n))$-bit float number precision, indicating that they are equivalent to constant-depth threshold circuits with polynomial size, and their ability is limited when handling more complex reasoning tasks which require higher-depth or nonuniform computations. As a result, Transformers with SMATs or AHATs are inherently unable to solve problems outside $\mathsf{TC}^{0}$, especially those that involve many inherently sequential computations. What about Transformers with CoT? Even though Transformers with CoT can address relatively more problems than CoT, Transformers still fail to solve problems requiring reasoning beyond $\mathsf{TC}^{0}$.
@@
SSMs have emerged as an alternative model to the popular LLMs, such as RNNs and Transformers. SSM presents ideal performance in tasks involving long-term dependencies and sequential reasoning [24]. The foundation of SSMs uses linear dynamical systems (LDS) or discrete-time state-space equations [24, 23] to represent the system’s internal state and its evolution over time. Using these mechanisms, SSMs are able to capture the sequential nature of data by updating the state iteratively, which has efficient inference and state-tracking [40, 22]. Compared to RNNs, SSMs have better scalability and stability when handling long sequences, and SSMs are capable of resolving the gradient-related issues inherent to RNNs [24].
@@ 3 Preliminaries
In Section 3.1, we introduce the float point number. In Section 3.2, we introduce the Mamba block.
@@ 3.1 Float Point Numbers
To compute SSM and Mamba correctly and effectively, we establish the computational framework by providing the definitions of the basic concepts of floating-point numbers and their related operations as follows.
@@
Consider $p\in\operatorname*{\mathbb{Z}}_{+}$. We have
@@ 3.2 Mamba Blocks
Now we introduce the softplus activation used in Mamba selection mechanisms as $\tau_{\Delta}$.
@@ 4.1 Approximating Logarithm in $\mathsf{TC}^{0}$
We use Taylor series expansion around 1 to approximate $\log(r)$, and we can get an approximation of $\log(r)$ with relative error bounded by $2^{-p-1}$.
We use the same technique, we can approximate $\log(2)$. Lastly, we compute $\log(x)$ as $\log(x)=\log(r)+k\cdot\log(2)$.
The $\mathsf{TC}^{0}$ circuit in constant depth can compute all operations. ∎
@@ 4.2 Recurrent SSMs are in $\mathsf{TC}^{0}$
In this section, we show recurrent SSM is in $\mathsf{TC}^{0}$.
@@
The computation of $\mathsf{SSM}_{\mathrm{recur}}(X)$ involves two primary steps: computing the hidden state updates $\mathcal{H}(X,A,B,\Delta)$ and iterative addition with multiplication. We can use a threshold circuit whose depth is
@@
Finally, we can show: $d_{\mathrm{recur}}=d_{h}+(d_{\mathrm{std}}+d_{\oplus}).$
Therefore, we get our desired result. ∎
@@ 4.3 Convolution SSMs are in $\mathsf{TC}^{0}$
In this section, we show convolution SSM is in $\mathsf{TC}^{0}$.
@@
whose depths are $d_{\mathrm{std}}+d_{\oplus}$ and $d_{\oplus}$, respectively.
Finally, we can conclude that: $d_{\mathrm{conv}}=d_{\mathrm{std}}+2d_{\oplus}$. Thus, we get the desired result. ∎
@@ 4.4 Circuit Complexity Bound for Selective SSM
whose depths are $d_{\mathrm{select}}$, $d_{\mathrm{disc}}$, $d_{\mathrm{recur}}$, and $d_{\mathrm{conv}}$ respectively.
Finally, we can show:
@@ 5 Hardness
To show Selective SSM and Mamba cannot solve arithmetic formula evaluation problems, Boolean formula value problems, and permutation composition problems. We leverage the difference between the complexity classes $\mathsf{TC}^{0}$ and $\mathsf{NC}^{1}$, under the assumption $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$.
Arithmetic formula evaluation problems, Boolean formula value problems, and permutation composition problems are defined to be $\mathsf{NC}^{1}$ problems in Section C.1C.2, and C.3.
From previous proof, we show Selective SSM and Mamba are both in $\mathsf{TC}^{0}$. Therefore, they cannot solve those $\mathsf{NC}^{1}$ problems. ∎
@@ 6 Conclusion
Our contributions include formal proofs of the circuit complexity bounds for Mamba and SSMs, and we show that their computational performances are equivalent to constant-depth uniform threshold circuits. Additionally, we provide hardness results. The hardness results show that these architectures cannot resolve sequential and state-dependent tasks that require higher computational depth. These new findings challenge the assumption that Mamba has higher computational capabilities than Transformers.
By building the theoretical limits of Mamba and SSMs, our work contributes to the broader understanding of the computational power of modern neural network models. We emphasize the need for future innovations to solve problems beyond $\mathsf{TC}^{0}$ so they can solve more complex and inherently sequential problems. We hope our study can inspire more research on designing newer architectures that can balance efficiency, scalability, and enhanced expressiveness.
@@ Appendix A Preliminaries
In this section, we introduce more definitions related to our work. In Section A.1, we introduce the circuit complexity classes. In Section A.2, we introduce more float point numbers and their operations. In Section A.3, we define the components of Recurrent SSM. In Section A.4, we define the components of Convolution SSM.
@@
A.1 Circuit Complexity
In this section, we provide an introduction to the fundamental concepts of circuit complexity classes.
We define the Boolean circuit below:
Definition A.1 (Boolean circuit, from Definition 6.1, On page 102 in [1]).
Definition A.2 (Circuit family recognizes languages, from Definition 6.2, On page 103 in [1]).
We now introduce $\mathsf{NC}^{i}$ class.
Definition A.3 ($\mathsf{NC}^{i}$ [1]).
$\mathsf{NC}^{i}$ consists of languages that can be decided by Boolean circuits with a size of $O(\text{poly}(n))$, depth $O((\log n)^{i})$, and utilizing ${\sf OR}$, ${\sf AND}$, and ${\sf NOT}$ gates with bounded fan-in.
Definition A.4 ($\mathsf{AC}^{i}$ [1]).
Definition A.5 ($\mathsf{TC}^{i}$ [67]).
Remark A.6.
In Definition A.5, ${\sf THRESHOLD}$ gates or ${\sf MOD}$ gates configured for prime values can replace ${\sf MAJORITY}$ gates. A Boolean circuit that includes any of these gates is referred to as a threshold circuit.
Definition A.7 ($\mathsf{P}$ [1]).
Fact A.8 (Hierarchy Folklore, [1], From Corollary 4.35, On page 110 in [1], in [67]).
Remark A.9.
Definition A.10 ($\mathsf{L}$-uniformity [1]).
$\mathsf{C}$ represents a language recognized by a circuit family $\mathcal{C}$, where $\mathcal{C}$ could be $\mathsf{NC}^{i}$, $\mathsf{AC}^{i}$, or $\mathsf{TC}^{i}$. Suppose we have a Turing machine that satisfying for any arbitrary $n\in\mathbb{N}$, computes a circuit in $\mathcal{C}$ for $n$ variables from the input $1^{n}$ using $O(\log n)$ space, such that the circuit $C_{n}$ recognizes $L$, then a language $L$, which is the subset of $\{0,1\}^{*}$, is said to be in $\mathsf{L}$-uniform $\mathsf{C}$.
Definition A.11 ($\mathsf{DLOGTIME}$-uniformity in [10]).
$\mathsf{C}$ is defined as in Definition A.10. Suppose we have a Turing machine that satisfying for any arbitrary $n\in\mathbb{N}$, computes $C_{n}$ in $\mathcal{C}$ for $n$ variables from the input $1^{n}$ within time $O(\log n)$, where $C_{n}$ recognizes $L$, then a language $L$, which is the subset of $\{0,1\}^{*}$, is said to be in $\mathsf{DLOGTIME}$-uniform $\mathsf{C}$.
A.2 Float Point Numbers
@@ C.4 Results About Hardness
It is by Theorem 4.4, Lemma C.3, and Fact A.8. ∎
@@
It is by Theorem 4.5, Lemma C.3, and Fact A.8. ∎
@@
It is by Theorem 4.4, Lemma C.7, and Fact A.8. ∎
@@
It is by Theorem 4.5, Lemma C.7, and Fact A.8. ∎
@@
It is by Theorem 4.4, Lemma C.12, and Fact A.8. ∎
@@
It is by Theorem 4.5, Lemma C.12, and Fact A.8. ∎
@@
It is by Theorem 4.4, Lemma C.13, and Fact A.8. ∎
@@
It is by Theorem 4.5, Lemma C.13, and Fact A.8. ∎

### v2

@@
Yifang Chen yifangc@uchicago.edu. The University of Chicago.    Xiaoyu Li xiaoyu.li2@student.unsw.edu.au. University of New South Wales.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at UC Berkeley.
@@ 1 Introduction
Sequential neural networks like RNNs, including their variants such as LSTMs and GRUs [49, 21], have good performance in capturing temporal dependencies and processing input step-by-step [18]. These advantages make them effective in tasks including time-series prediction [7] and speech recognition [104]. Traditional RNNs [50] and their enhanced variance, LSTMs perform well in testing because of their sequential nature, but their training times tend to be slow and suffer from vanishing or exploding gradient issues, which limit their capabilities to capture long-term dependencies [122]. Transformers [117], equipped with a self-attention mechanism, provides an efficient solution to the slow training problem by enabling parallelized computations. Large Language Models (LLMs) based on the Transformer architecture, such as GPT-4 [99], GPT-4o [100], OpenAI’s o1 [101], Llama 3.1 [94], Claude [6], and Gemini [39], have become ubiquitous nowadays, and their integrations into modern technology reshaped our expectations of the limits of their capabilities. Transformers are capable of training efficiently on large datasets, but their quadratic memory and time complexity with respect to sequence length make them expensive in resources, both in terms of memory and processing power, during training and inference. Specifically, self-attention mechanisms grows $O(n^{2})$ in terms of computational complexity [69].
@@
Section 2 introduces the works related to our paper. Section 3 introduces key computational concepts and Mamba definitions that form the basis for subsequent sections. Then, we present the circuit complexity results for Selective SSM and Mamba in Section 4. Section 5 details our hardness results. Finally, Section 6 gives a conclusion.
@@ 2 Related Work
Circuit Complexity, a crucial set of metrics in computational complexity theory, studies the computational power of circuit families. It has valuable applications in comprehending the capabilities of machine learning models [102, 47, 46, 97, 96, 34, 80, 125, 28, 128, 64, 22]. The complexity classes include $\mathsf{AC^{0}}$ represents problems that are highly parallelizable equipped with standard logic gates, which can be solved by constant-depth circuits with unbounded fan-in $\mathsf{AND}$, $\mathsf{OR}$, and $\mathsf{NOT}$ gates; $\mathsf{TC^{0}}$ class extends from $\mathsf{AC^{0}}$ with additional majority gates; $\mathsf{NC^{1}}$ problems can be solved by $O(\log n)$-depth circuits with bounded fan-in. These circuit complexity classes form a hierarchy: $\mathsf{AC}^{0}\subset\mathsf{TC}^{0}\subseteq\mathsf{NC}^{1}$ [97]. The question of whether $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$ remains an open topic of discussion. [63] demonstrates that while Transformers can simulate nonsolvable semi-automata, their depth is influenced by the length of the input sequence. Building on this, [80] investigates the expressive power of Transformers augmented with Chain-of-Thought (CoT) reasoning in the context of circuit complexity. They propose the following relationships:
@@
Transformers have shown outstanding performance on tasks from natural language processing, but they present limited effectiveness in mathematical computations. A series of research highlights the reasoning limitations of Transformer Model [3, 96, 17, 120, 84, 19, 61][19] shows that average-hard attention transformers (AHATs) and softmax-attention transformers (SMATs) are in $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ with $O(\operatorname{poly}(n))$-bit float number precision, indicating that they are equivalent to constant-depth threshold circuits with polynomial size, and their ability is limited when handling more complex reasoning tasks which require higher-depth or nonuniform computations. As a result, Transformers with SMATs or AHATs are inherently unable to solve problems outside $\mathsf{TC}^{0}$, especially those that involve many inherently sequential computations. What about Transformers with CoT? Even though Transformers with CoT can address relatively more problems than CoT, Transformers still fail to solve problems requiring reasoning beyond $\mathsf{TC}^{0}$.
@@
SSMs have emerged as an alternative model to the popular LLMs, such as RNNs and Transformers. SSM presents ideal performance in tasks involving long-term dependencies and sequential reasoning [37]. The foundation of SSMs uses linear dynamical systems (LDS) or discrete-time state-space equations [37, 36] to represent the system’s internal state and its evolution over time. Using these mechanisms, SSMs are able to capture the sequential nature of data by updating the state iteratively, which has efficient inference and state-tracking [58, 35]. Compared to RNNs, SSMs have better scalability and stability when handling long sequences, and SSMs are capable of resolving the gradient-related issues inherent to RNNs [37] and have recently garnered attention for their versatility across various tasks such as sequential recommendation [76, 65] and image deblurring [78].
@@ 3 Preliminaries
In Section 3.1, we introduce the circuit complexity classes. In Section 3.2, we introduce the float point number. In Section 3.3, we introduce the Mamba block.
@@
3.1 Circuit Complexity
In this section, we provide an introduction to the fundamental concepts of circuit complexity classes. We define the Boolean circuit below:
Definition 3.1 (Boolean circuit, from Definition 6.1, On page 102 in [2]).
Definition 3.2 (Circuit family recognizes languages, from Definition 6.2, On page 103 in [2]).
We now introduce $\mathsf{NC}^{i}$ class.
Definition 3.3 ($\mathsf{NC}^{i}$ [2]).
$\mathsf{NC}^{i}$ consists of languages that can be decided by Boolean circuits with a size of $O(\operatorname{poly}(n))$, depth $O((\log n)^{i})$, and utilizing ${\sf OR}$, ${\sf AND}$, and ${\sf NOT}$ gates with bounded fan-in.
Definition 3.4 ($\mathsf{AC}^{i}$ [2]).
Definition 3.5 ($\mathsf{TC}^{i}$ [118]).
Remark 3.6.
In Definition 3.5, ${\sf THRESHOLD}$ gates or ${\sf MOD}$ gates configured for prime values can replace ${\sf MAJORITY}$ gates. A Boolean circuit that includes any of these gates is referred to as a threshold circuit.
Definition 3.7 ($\mathsf{P}$ [2]).
Fact 3.8 (Hierarchy Folklore, [2], From Corollary 4.35, On page 110 in [2], in [118]).
Remark 3.9.
Definition 3.10 ($\mathsf{L}$-uniformity [2]).
$\mathsf{C}$ represents a language recognized by a circuit family $\mathcal{C}$, where $\mathcal{C}$ could be $\mathsf{NC}^{i}$, $\mathsf{AC}^{i}$, or $\mathsf{TC}^{i}$. Suppose we have a Turing machine that is satisfying for any arbitrary $n\in\mathbb{N}$, computes a circuit in $\mathcal{C}$ for $n$ variables from the input $1^{n}$ using $O(\log n)$ space, such that the circuit $C_{n}$ recognizes $L$, then a language $L$, which is the subset of $\{0,1\}^{*}$, is said to be in $\mathsf{L}$-uniform $\mathsf{C}$.
Definition 3.11 ($\mathsf{DLOGTIME}$-uniformity in [14]).
$\mathsf{C}$ is defined as in Definition 3.10. Suppose we have a Turing machine that satisfying for any arbitrary $n\in\mathbb{N}$, computes $C_{n}$ in $\mathcal{C}$ for $n$ variables from the input $1^{n}$ within time $O(\log n)$, where $C_{n}$ recognizes $L$, then a language $L$, which is the subset of $\{0,1\}^{*}$, is said to be in $\mathsf{DLOGTIME}$-uniform $\mathsf{C}$.
@@ 3.2 Float Point Numbers
To compute SSM and Mamba correctly and effectively, we establish the computational framework by providing the definitions of the basic concepts of floating-point numbers and their related operations.
@@
Let $p\in\operatorname*{\mathbb{Z}}_{+}$. We have
@@ 3.3 Mamba Blocks
Now, we introduce the softplus activation used in Mamba selection mechanisms as $\tau_{\Delta}$.
@@ 4.1 Approximating Logarithm in $\mathsf{TC}^{0}$
We use Taylor series expansion around 1 to approximate $\log(r)$, and we can get an approximation of $\log(r)$ with relative error bounded by $2^{-p-1}$. Using the same technique, we can approximate $\log(2)$. Lastly, we compute $\log(x)$ as $\log(x)=\log(r)+k\cdot\log(2)$. The $\mathsf{TC}^{0}$ circuit in constant depth can compute all operations. ∎
@@ 4.2 Recurrent SSMs are in $\mathsf{TC}^{0}$
In this section, we show recurrent SSM is in $\mathsf{TC}^{0}$. We provide more details about recurrent SSM in Appendix A.2.
@@
The computation of $\mathsf{SSM}_{\mathrm{recur}}(X)$ involves two primary steps: computing the hidden state updates $\mathcal{H}(X,A,B,\Delta)$ and iterative addition with multiplication. We use a threshold circuit whose depth is
@@
Finally, we can show: $d_{\mathrm{recur}}=d_{h}+(d_{\mathrm{std}}+d_{\oplus}).$ Therefore, we get our desired result. ∎
@@ 4.3 Convolution SSMs are in $\mathsf{TC}^{0}$
In this section, we show convolution SSM is in $\mathsf{TC}^{0}$. We provide more details about recurrent SSM in Appendix A.3.
@@
whose depths are $d_{\mathrm{std}}+d_{\oplus}$ and $d_{\oplus}$, respectively. Finally, we can conclude that: $d_{\mathrm{conv}}=d_{\mathrm{std}}+2d_{\oplus}$. Thus, we get the desired result. ∎
@@ 4.4 Circuit Complexity Bound for Selective SSM
whose depths are $d_{\mathrm{select}}$, $d_{\mathrm{disc}}$, $d_{\mathrm{recur}}$, and $d_{\mathrm{conv}}$ respectively. Finally, we can show:
@@ 4.5 Circuit Complexity Bound for Mamba
Theorem 4.5 demonstrates that a $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ circuit family can simulate Mamba, showing the Mamba representation capacity limitations. In previous work, [95] showed that SSM and Mamba can be simulated by $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ with $c\log(n)$ precision. However, we improve the uniformity and precision in [95] by proving that Mamba can be simulated by $\mathsf{DLOGTIME}$-uniform $\mathsf{TC}^{0}$ with $\mathrm{poly}(n)$ precision by new techniques introduced from [19]. Our complexity bound is better than previous work.
@@ 5 Hardness
To show Selective SSM and Mamba cannot solve arithmetic formula evaluation problems, Boolean formula value problems, and permutation composition problems. We leverage the difference between the complexity classes $\mathsf{TC}^{0}$ and $\mathsf{NC}^{1}$, under the assumption $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$. Arithmetic formula evaluation problems, Boolean formula value problems, and permutation composition problems are defined to be $\mathsf{NC}^{1}$ problems in Section C.1C.2, and C.3. From previous proof, we show Selective SSM and Mamba are both in $\mathsf{TC}^{0}$. Therefore, they cannot solve those $\mathsf{NC}^{1}$ problems. ∎
To the best of our knowledge, there is no previous work proving that Mamba and SSM with $\mathrm{poly}(n)$ precision cannot solve arithmetic formula problems, boolean formula value problems, and permutation composition problems.
@@ 6 Conclusion
Our contributions include formal proofs of the circuit complexity bounds for Mamba and SSMs, and we show that their computational performances are equivalent to constant-depth uniform threshold circuits. Additionally, we provide hardness results. The hardness results show that these architectures cannot resolve sequential and state-dependent tasks that require higher computational depth. These new findings challenge the assumption that Mamba has higher computational capabilities than Transformers. By building the theoretical limits of Mamba and SSMs, our work contributes to the broader understanding of the computational power of modern neural network models. We emphasize the need for future innovations to solve problems beyond $\mathsf{TC}^{0}$ so they can solve more complex and inherently sequential problems. We hope our study can inspire more research on designing newer architectures that can balance efficiency, scalability, and enhanced expressiveness.
@@ Appendix A Preliminaries
In this section, we introduce more definitions related to our work. In Section A.1, we introduce more float point numbers and their operations. In Section A.2, we define the components of Recurrent SSM. In Section A.3, we define the components of Convolution SSM.
@@
A.1 Float Point Numbers
@@ C.4 Results About Hardness
It is by Theorem 4.4, Lemma C.3, and Fact 3.8. ∎
@@
It is by Theorem 4.5, Lemma C.3, and Fact 3.8. ∎
@@
It is by Theorem 4.4, Lemma C.7, and Fact 3.8. ∎
@@
It is by Theorem 4.5, Lemma C.7, and Fact 3.8. ∎
@@
It is by Theorem 4.4, Lemma C.12, and Fact 3.8. ∎
@@
It is by Theorem 4.5, Lemma C.12, and Fact 3.8. ∎
@@
It is by Theorem 4.4, Lemma C.13, and Fact 3.8. ∎
@@
It is by Theorem 4.5, Lemma C.13, and Fact 3.8. ∎